Ex 18-7 Equivalen units of production The following information concerns production in the Baking Department for May. All dirPlease answer Ex 18-8. Thanks
Answer #1
Baking Department
FIFO Method:
Equivalent Units
DM Conversion costs
Whole units % Units % Units
Opening WIP 6500
Started 135000
Total Units to account for 141500
Opening WIP 6500 0 0 40 2600
Completed and tfd.( 137200 - 6500) 130700 100 130700 100 130700
Closing WIP 4300 100 4300 60 2580
Total Units to account for ( EUP) 141500 135000 135880
Cost Information
Costs to account for: Materials Conversion Total
Op. WIP 12480 12480
Added during the period 229500 82308 311808 ( 62300 + 20008 = 82308 )
Total costs to account for 229500 94788 324288
Cost per EUP :-
Cost to be accounted for 229500 82308 311808
EUP 135000 135880
Cost per EUP 1.7 0.605740359 2.305740359
Costs Accounted for: Materials Conversion Total
Completed and tfd.
Opening WIP 0 12480 12480
Completed and tfd. 130700 * 1.7 222190 133300 * 0.606 80745.18987 302935.1899
Completed and tfd. 222190 93225.18987 315415.1899
Cl. WIP 4300 * 1.7 7310 2580 * 0.606 1562.810127 8872.810127
Total costs accounted for 229500 94788 324288
Cost reconciliation schedule
Costs Accounted for:
Transferred out 315415.1899
closing WIP
Materials 7310
Conversion costs 1562.810127 8872.810127
Total costs 324288
Since the format was not available, the question has been solved using this format.
Required answers :-
a.1 - $ 1.7
a.2 - $ 0.606
a.3 - 2600 * 0.606 = $ 1576 +12480 = $ 14056 ( rounding off difference can be )
a.4 - 130700 * 2.306 = 301394
a.5 - $ 8873
b. CC per EUP ( opening WIP ) =( 12480 - 6500 * 1.7 )/ 6500 * 60 % = 1430 / 3900 = $ 0.37
It increased.